What is the boiling point of a solution of cane sugar, C12H22O11, containing 100 g of the solute per kilogram of water?


1
Expert's answer
2021-03-08T06:13:48-0500

Solution:

A solution will boil at a higher temperature than the pure solvent. This is the colligative property called boiling point elevation.

An equation has been developed for this behavior. It is:

Δt = i × Kb × m

Kb for water is 0.52°C kg mol¯1

i = 1 (for nonelectrolyte - C12H22O11)

m = the molality of the solute


Molality of C12H22O11 = Moles of C12H22O11 / kilograms solvent

Moles of C12H22O11 = Mass of C12H22O11 / Molar mass of C12H22O11

The molar mass of C12H22O11 is 342.3 g mol-1.

Hence,

Moles of C12H22O11 = (100 g) / (342.3 g mol-1) = 0.292 mol

Molality of C12H22O11 = (0.292 mol) / (1 kg) = 0.292 mol/kg


Thus:

Δt = i × Kb × m = (1) × (0.52°C kg mol¯1) × (0.292 mol/kg) = 0.15184°C = 0.15°C


The boiling point of pure water is 100°C.

Hence,

t = 100°C + 0.15°C = 100.15°C

t = 100.15°C


Answer: The boiling point of a solution of cane sugar (C12H22O11) is 100.15°C.